# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v1

@@ B.2 Proof of Lemma 3
From Lemmas 17 and 18, we obtain the following corollary:
@@ C.2 Proof of Lemma 7
$\Gamma=I-Z^{\dagger}Z\geq P-Z^{\dagger}Z=P^{2}-Z^{\dagger}Z=(C^{+})^{\dagger}(C^{\dagger}C-C^{\prime{\dagger}}C^{\prime})C^{+}\geq(C^{+})^{\dagger}C^{+}\geq 0,$ (221)
@@
From the definition of $G_{\mathrm{succ.}}(\delta)$, we have
$G_{\mathrm{succ.}}(\delta)=\bigcup_{n\in\mathbb{N}}H\left(\delta-\frac{1}{n}\right).$ (256)
Therefore, $G_{\mathrm{succ.}}(\delta)$ is a countable union of closed sets and hence is a measurable set.
@@ C.4 Proof of Lemma 8
Let $G$ be a compact Lie group and $\rho$ be an arbitrary state. Fix $\epsilon\in(0,1/2)$. Then there exists an estimator of $g\in G$ by using consuming $\mathcal{U}_{g}(\rho)^{\otimes n}$ with $n=N^{1-\epsilon}$ such that its success probability satisfies
$\lim_{N\to\infty}\mathrm{Pr}_{\mathcal{U}_{g}(\psi)^{\otimes N^{1-\epsilon}}}(\|\bm{\theta}\|<N^{-1/2+\epsilon})=1.$ (304)
@@
by performing measurements on $n=N^{1-\epsilon}$ copies of the system in a state $\mathcal{U}_{g}(\rho)^{\otimes n}$. From the estimated state $\hat{\rho}_{n}$, we pick up any $\hat{g}\in G$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$ as an estimate of the true value $g\in G$. If there is no $\hat{g}$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$, which may happen with an exponentially small probability, we arbitrarily select $\hat{g}$ as an element in $G$, such as $e\in G$. Whenever $\|\hat{\rho}_{n}-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/(2c)$ holds, the estimated value $\hat{g}\in G$ satisfies $\|\mathcal{U}_{\hat{g}}(\rho)-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/c$ due to the triangle inequality. In this case, for all sufficiently large $N$ such that $N^{-1/2+\epsilon}/c<\delta_{*}$, we get $\|\bm{\theta}\|<N^{-1/2+\epsilon}$. Therefore, the success probability is bounded as

### v2

@@ B.3 Proof of Lemma 4
From Lemmas 22 and 23, we obtain the following corollary:
@@ C.2 Proof of Lemma 8
$\Gamma=I-Z^{\dagger}Z\geq P-Z^{\dagger}Z=P^{2}-Z^{\dagger}Z=(C^{+})^{\dagger}(C^{\dagger}C-C^{\prime{\dagger}}C^{\prime})C^{+}\geq 0,$ (258)
@@
From the definition of $G_{\mathrm{succ.}}^{(g,\delta)}$, we have
$G_{\mathrm{succ.}}^{(g,\delta)}=\bigcup_{n\in\mathbb{N}}H\left(\delta-\frac{1}{n}\right).$ (293)
Therefore, $G_{\mathrm{succ.}}^{(g,\delta)}$ is a countable union of closed sets and hence is a measurable set.
@@ C.4 Proof of Lemma 9
Let $G$ be a compact Lie group and $\rho$ be an arbitrary state. Fix $\epsilon\in(0,1/2)$. Then there exists an estimator of $g\in G$, which consumes $\mathcal{U}_{g}(\rho)^{\otimes n}$ with $n=\lceil N^{1-\epsilon}\rceil$, such that its worst-case success probability satisfies
$\lim_{N\to\infty}p^{\mathrm{succ.}}(N^{-1/2+\epsilon})=1.$ (341)
@@
by performing measurements on $n=\lceil N^{1-\epsilon}\rceil$ copies of the system in a state $\mathcal{U}_{g}(\rho)^{\otimes n}$, where we have used $nN^{-1+2\epsilon}\geq N^{1-\epsilon}N^{-1+2\epsilon}=N^{\epsilon}$. From the estimated state $\hat{\rho}_{n}$, we pick up any $\hat{g}\in G$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$ as an estimate of the true value $g\in G$. If there is no $\hat{g}$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$, which may happen with an exponentially small probability, we arbitrarily select $\hat{g}$ as an element in $G$, such as $e\in G$. Whenever $\|\hat{\rho}_{n}-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/(2c)$ holds, the estimated value $\hat{g}\in G$ satisfies $\|\mathcal{U}_{\hat{g}}(\rho)-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/c$ due to the triangle inequality. In this case, for all sufficiently large $N$ such that $N^{-1/2+\epsilon}/c<\delta_{*}$, we get $\|\bm{\theta}\|<N^{-1/2+\epsilon}$. Therefore, the success probability is bounded as
$\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes n})\chi_{G_{\mathrm{succ.}}^{(g,\delta)}}(\hat{g})\geq 1-c_{1}e^{-c_{2}N^{\epsilon}},$ (343)
where $c_{1}\coloneqq d$ and $c_{2}\coloneqq\frac{1}{172g(d)r^{2}c^{2}}$. In other words, the failure probability is exponentially small in $N$. Since the rank of the state, $r$, is invariant under unitary transformation and hence the constants $c_{1},c_{2}$ are independent of $g$, this bound is uniform in $g\in G$, implying that
$p^{\mathrm{succ.}}(N^{-1/2+\epsilon})=\inf_{g\in G}\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes n})\chi_{G_{\mathrm{succ.}}^{(g)}(\delta)}(\hat{g})\geq 1-c_{1}e^{-c_{2}N^{\epsilon}}.$ (344)
Therefore, we get
$\lim_{N\to\infty}p^{\mathrm{succ.}}(N^{-1/2+\epsilon})=1.$ (345)
Appendix D Proof of Lemma 10
We here prove Lemma 10 by using Lemma 8 and the relation in Eq. (85).
Lemma 38 (Restatement of Lemma 10).
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let |$\ket{\psi}\in\mathcal{H}$ and |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ be pure states. We define $\psi_{g}\coloneqq\mathcal{U}_{g}(\psi)$ and $\phi_{g}\coloneqq\mathcal{U}^{\prime}_{g}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi_{g}}$ and $\mathcal{Q}^{\phi_{g}}$. If $r>0$ satisfies $\mathcal{Q}^{\psi_{g}}\geq r\mathcal{Q}^{\phi_{g}}$ for all $g\in G$, then, there exists a sequence of quantum channels $\{\mathcal{E}^{(g)}_{N}\}_{N}$ such that the conversion error
$\delta_{N}(g,\bm{u})\coloneqq T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\phi_{g}\right)^{\otimes\lfloor rN\rfloor}\right)$ (346)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}(g,\bm{u})=0$ for $\epsilon\in(0,1/9)$.
Proof.
Let $G=\bigsqcup_{i=0}^{k}G_{i}$ be the decomposition of a compact Lie group $G$ into its connected components. We arbitrarily fix $g_{i}\in G_{i}$ for $i=0,\cdots,k$. Applying Lemma 8 to $\psi_{g_{i}}\coloneqq\mathcal{U}_{g_{i}}$ and $\phi_{g_{i}}\coloneqq\mathcal{U}^{\prime}_{g_{i}}(\phi)$ for each $g_{i}$, we find that there exists a sequence of quantum channels $\{\mathcal{E}_{N}^{(g_{i})}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi_{g_{i}})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi_{g_{i}})^{\otimes\lfloor rN\rfloor}\right)=0$ (347)
for $\epsilon\in(0,1/9)$.
For a general $g\in G$, if $g_{i}$ belongs to a connected component $g\in G_{i}$, we define a channel $\mathcal{E}_{N}^{(g)}$ by using $\mathcal{E}_{N}^{(g_{i})}$ as follows: Let $g^{\prime}\in G_{0}$ be an element in the identity component satisfying $g=hg_{i}$. From Eq. (85), there exists an invertible $\dim G\times\dim G$ real matrix $V(h)$ satisfying
$U(h)^{\dagger}X_{\mu}U(h)=\sum_{\nu=1}^{\dim G}V(h)_{\nu\mu}X_{\nu}.$ (348)
We define $\mathcal{E}_{N}^{(g)}\coloneqq\left(\mathcal{U}^{\prime}_{h}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}^{(g_{i})}\circ\left(\mathcal{U}_{h}^{-1}\right)^{\otimes N}$. Since $U(h)^{\dagger}(\bm{u}\cdot\bm{X})U(h)=\bm{u}\cdot V(h)^{\top}\bm{X}=\bm{v}\cdot\bm{X}$, where $\bm{v}\coloneqq\bm{u}V(h)^{\top}$, we have
$\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right)$ $=\left(\mathcal{U}_{h}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{h}^{-1}\circ\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\circ\mathcal{U}_{h}\left(\psi_{g_{i}}\right)\right)^{\otimes N}$ (349)
$=\left(\mathcal{U}^{\prime}_{h}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)\right).$ (350)
Therefore, we get
$T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)$ $=T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right),$ (351)
where we have used $U^{\prime}(h)^{\dagger}(\bm{u}\cdot\bm{X}^{\prime})U^{\prime}(h)=\bm{v}\cdot\bm{X}^{\prime}$. Note that the induced norm $\|V(h)\|_{\infty}\coloneqq\max_{\bm{u}\neq\bm{0}}\|V(h)\bm{u}\|/\|\bm{u}\|$ is finite for any $h\in G_{0}$. Since $\|V(h)\|_{\infty}$ and continuous in $h\in G_{0}$ and $G_{0}$ is compact, we find $c\coloneqq\sup_{h\in G_{0}}\|V(h)\|_{\infty}$ is also finite. By using $\|\bm{v}\|\leq c\|\bm{u}\|$, for any $g\in G$ and $\|\bm{u}\|<N^{\epsilon}$, we get
$\delta_{N}(g,\bm{u})$ $\coloneqq T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)$ (352)
$\leq\max_{i\in\{0,\cdots,k\}}\sup_{\|\bm{v}\|<cN^{\epsilon}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right).$ (353)
Since $c$ is independent of $N$, for any $\epsilon\in(0,1/9)$, $cN^{\epsilon}<N^{\epsilon^{\prime}}$ holds for all sufficiently large $N$ for any $\epsilon^{\prime}\in(\epsilon,1/9)$. Therefore,
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}(g,\bm{u})$ $\leq\lim_{N\to\infty}\max_{i\in\{0,\cdots,k\}}\sup_{\|\bm{v}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right)$ (354)
$=\max_{i\in\{0,\cdots,k\}}\lim_{N\to\infty}\sup_{\|\bm{v}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right)$ (355)
$=0,$ (356)
where the last equality follows from Eq. (347). ∎
Appendix E Proof of Eq. (75)
Following the standard argument using typical sequence [72, 21], we prove Eq. (75), i.e., $\sum_{i\in[k]}p_{i}\mathcal{A}_{\mathrm{c}}(\rho_{i})\geq\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho=\sum_{i\in[k]}p_{i}\rho_{i}$, where $[k]\coloneqq\{1,2,\cdots,k\}$.
Proof.
Without loss of generality, we can assume that $p_{i}>0$ for all $i\in[k]$. If $\mathcal{A}_{\mathrm{c}}(\rho_{i})=\infty$, the inequality is trivial. Therefore, we only consider the case where $\mathcal{A}_{\mathrm{c}}(\rho_{i})<\infty$ for all $i\in[k]$ below.
The i.i.d. copies of $\rho$ is given by
$\rho^{\otimes N}=\left(\sum_{i\in[k]}p_{i}\rho_{i}\right)^{\otimes N}=\sum_{\bm{i}\in[k]^{N}}p_{\bm{i}}\rho_{\bm{i}},$ (357)
where $\bm{i}\coloneqq i_{1}i_{2}\cdots i_{N}\in[k]^{N}$, $p_{\bm{i}}\coloneqq\prod_{j=1}^{N}p_{i_{j}}$, and $\rho_{\bm{i}}\coloneqq\bigotimes_{j=1}^{N}\rho_{i_{j}}$. For $l\in[k]$, let $n_{l}(\bm{i})$ denote the number of occurrence of $l$ in $i_{1}i_{2}\cdots i_{N}$. For $\delta>0$, we define $\delta$-typical sequence as
$\mathcal{T}_{\delta}\coloneqq\left\{\bm{i}=i_{1}i_{2}\cdots i_{N}\in[k]\,\middle|\,\forall l\in[k],\,\left|\frac{n_{l}(\bm{i})}{N}-p_{l}\right|\leq\delta\right\}.$ (358)
The i.i.d. state $\rho^{\otimes N}$ is decomposed as
$\rho^{\otimes N}=\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\rho_{\bm{i}}+\sum_{\bm{i}\notin\mathcal{T}_{\delta}}p_{\bm{i}}\rho_{\bm{i}}.$ (359)
Let us introduce a density operator
$\tilde{\rho}_{N}\coloneqq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\rho_{\bm{i}}+p_{\mathrm{err}}\rho_{\mathrm{sym}},$ (360)
where $p_{\mathrm{err}}\coloneqq 1-\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}$ and $\rho_{\mathrm{sym}}$ is some symmetric state, e.g., the maximally mixed state. From the standard argument on a typical sequence, for any $\delta>0$, the probability $p_{\mathrm{err}}$ converges to zero as $N\to\infty$, which implies $\lim_{N\to\infty}T\left(\rho^{\otimes N},\tilde{\rho}_{N}\right)=0$. Thus, we prepare $\tilde{\rho}_{N}$ instead of $\rho^{\otimes N}$.
The state $\rho_{\bm{i}}$ is equivalent to $\bigotimes_{l\in[k]}\rho_{l}^{\otimes n_{l}(\bm{i})}$ up to permutation. Note that the permutation, which is composed of the SWAP operations among subsystems, is a $G$-covariant operation. For a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$, $n_{l}(\bm{l})\leq N(p_{l}+\delta)$ holds for all $l\in[k]$. Therefore, a state $\rho_{\bm{i}}$ for a typical sequence $\bm{i}\in\mathcal{T}_{\delta}$ can be created by first preparing $\bigotimes_{l\in[k]}\rho_{l}^{\otimes\lfloor N(p_{l}+\delta)\rfloor}$ via a $G$-covariant operation, discarding some copies if needed, and then operating a permutation. For each $l\in[k]$, for $r_{l}$ satisfying $r_{l}>(p_{l}+\delta)\mathcal{A}_{\mathrm{c}}(\rho_{l})$, it holds
$\{\phi^{\otimes\lceil r_{l}N\rceil}\}_{N}\xrightarrow{G\text{-cov.}}\{\rho_{l}^{\otimes\lfloor N(p_{l}+\delta)\rfloor}\}_{N}.$ (361)
Therefore, introducing $r\coloneqq\sum_{i\in[k]}r_{l}$, there exists a sequence of $G$-covariant channels $\{\mathcal{E}_{N}^{(\bm{i})}\}_{N}$ for any $\bm{i}\in\mathcal{T}_{\delta}$ such that
$\lim_{N\to\infty}T\left(\mathcal{E}_{N}^{(\bm{i})}(\phi^{\otimes\lceil rN\rceil}),\rho_{\bm{i}}\right)=0.$ (362)
Defining a sequence of $G$-covariant channels $\{\mathcal{E}_{N}\}_{N}$ by
$\mathcal{E}_{N}(\cdot)\coloneqq\sum_{\bm{i}\in\mathcal{T}_{\delta}}p_{\bm{i}}\mathcal{E}_{N}^{(\bm{i})}(\cdot)+p_{\mathrm{err}}\mathcal{E}_{\mathrm{sym},N}(\cdot),$ (363)
where $\mathcal{E}_{\mathrm{sym},N}(\cdot)$ denotes the $G$-covariant channel that maps arbitrary state into a symmetric state $\rho_{\mathrm{sym}}$, we get